OpenStudy (anonymous):

Factor the trinomial, if possible. 8a^2+11ab-10b^2

OpenStudy (anonymous):

ac=8*-10=-80 you want 2 factors of -80 that add up to 11, both have different signs 16 and -5 8a^2-16ab+5ab-10b^2= 8a(a-2b)+5b(a-2b)= (a-2b)(8a+5b)

OpenStudy (anonymous):

I now see though, that you missplaced the signs for the 16ab and the -5ab in your equation. It should be (a+2b)(8a-5b). Thanks for walking me through the process.

OpenStudy (anonymous):

ah you're right, i did misplace the signs lol